OpenStudy (anonymous):

So i have a problem and it's about inequality, and it required me to answer it in a complete sentence, little help please :) Problem: Glenda has a $40 gift certificate to a cafe that sells her favorite tuna sandwich for $3.75 after tax. What are the possible number of tuna sandwiches that Glenda can buy with her gift certificate?! And it even said to define a variable, write an inequality then solve. Anyone help me? Thanks :)

OpenStudy (jusaquikie):

3.75x<40 x<40/3.75 let me know if you understand, i'll hang out a bit to answer your questions.

OpenStudy (anonymous):

Kk so explanation please :) I dont really get it

OpenStudy (jusaquikie):

is this for algebra?

OpenStudy (anonymous):

pre-algerba actually

OpenStudy (jusaquikie):

do you know how to solve for x? as in 3x=15

OpenStudy (anonymous):

you minus it?

OpenStudy (anonymous):

-3 on each side?

OpenStudy (jusaquikie):

ok think of the equal sign as line, and you want to balance each side out

OpenStudy (jusaquikie):

if you have 3 times some number = 15 then 3 times what would give you 15?

OpenStudy (anonymous):

kk keep it going :)

OpenStudy (jusaquikie):

x is just a number that we don't know what it is yet but we knw that 3*(this number) is 15

OpenStudy (jusaquikie):

now if we didn't know that 3*5 is 15 we could take the 3 to the other side and then all we have on the 1 side is x=

OpenStudy (anonymous):

12?

OpenStudy (jusaquikie):

that would be right it it were 3+x=15 but if it's 3*x we have to divide 3

OpenStudy (anonymous):

oops nvm i get it

OpenStudy (anonymous):

ohh i was right nvm again :P

OpenStudy (jusaquikie):

think of the equal sign like the opposite side, so if you 3*5=15 then 3=15/5

OpenStudy (jusaquikie):

if it's multiplied on 1 side then to move it you devide the other side by what your moving, if it's just + or - then you change the sign and add or subtract from the other side

OpenStudy (anonymous):

so 3x=15 then its gonna be 15+3=x?

OpenStudy (anonymous):

can you put it in a simple way ?

OpenStudy (jusaquikie):

no if x is next to a number without a + or minus between that means x multiplied by the number

OpenStudy (jusaquikie):

3x means 3 times x

OpenStudy (anonymous):

ok gotcha i know that but how to make it balance on either sides?

OpenStudy (jusaquikie):

just like x/3 is x divided by 3

OpenStudy (anonymous):

so x=15?

OpenStudy (anonymous):

because 15/3=3

OpenStudy (jusaquikie):

so if you had \[\frac{ x }{ 3 }=5\] that would mean that you would need to multiply both sides by 3 so you would get \[\frac{ 3x }{ 3 }=5*3\] the 3 and 3 cancle out and your left with x=3*5 or 15

OpenStudy (jusaquikie):

yes i think your starting to understand a bit

OpenStudy (jusaquikie):

they are still equal if you do the same thing to both sides so to move 3+x=15 we would subtract 3 from both sides and we get x+12 if we have 3x=15 we have to divide both sides by 3 so we get x=15/3 or 5 if we have x/3= 5 we multiply both sides by 3 and get x= 3*5 or 15

OpenStudy (anonymous):

OH i get it so back to the problem 3x=15 would be x=5?

OpenStudy (jusaquikie):

< and > work the same way if 3+2>4 then 3>4-2

OpenStudy (anonymous):

oh ok get it:)

OpenStudy (jusaquikie):

yes that is correct lets' try this one to be sure you understand \[4x=20\] what is x?

OpenStudy (anonymous):

So what about this prbolem? : 2/3 > 1/2 - d/6 / <---- mean " over". Like 2 over 3, 1 over 2 and d over 6

OpenStudy (jusaquikie):

ok treat the > just like you would = what would you do first?

OpenStudy (jusaquikie):

you are trying to get d by itself

OpenStudy (anonymous):

kick the d out?

OpenStudy (jusaquikie):

no we want to save that for last wet's get the addition/subtraction out of the way first, \[\frac{ 2 }{3 }>\frac{ 1 }{ 2 }-\frac{ d }{ 6 }\] so if we move 1/2 over it becomes -1/2

OpenStudy (anonymous):

how?

OpenStudy (jusaquikie):

so\[\frac{ 2 }{ 3 }-\frac{ 1 }{ 2 }>\frac{ d }{ 6 }\] now how do we subtract fractions?

OpenStudy (anonymous):

uhmm divide it?

OpenStudy (anonymous):

OMG im so dump. SOrry im just wasting your time :P

OpenStudy (jusaquikie):

you need a common denominator, nah it's ok i need a break from my Calculus homework lol

OpenStudy (anonymous):

Oh lol you doing a 7th grader math hw

OpenStudy (jusaquikie):

don't worry about the d/6 right now one piece at a time, math is all in baby steps, when we try to skip stuff we make mistakes so think of it as doing what we know how to make it simpler and right now we can make it simple by subtracting our fractions 2/3-1/2

OpenStudy (jusaquikie):

yeah some day you can help out a seventh grader lol and it's all be good

OpenStudy (anonymous):

haha thanks :) OK back to the prob. so 2/3 - 1/2 gonna be 1?

OpenStudy (jusaquikie):

so to subtract the denominator has to be the same so we make them the same by finding something in common or cross multiplying so \[\frac{ 2*2 }{ 3*2 } - \frac{ 1*3 }{ 2*3 } = \frac{ 4 }{ 6 }-\frac{ 3 }{6 }\]

OpenStudy (anonymous):

oh yeah i forgot im not that good at math :P or sp its gonna be 1/6 ?

OpenStudy (jusaquikie):

now we have \[\frac{ 1 }{ 6 }>\frac{ d }{6 }\] what would happen if we multiplied both sides by 6?

OpenStudy (anonymous):

uhm it still gonna be the same number. But since it lready 6 so i dont think we have to

OpenStudy (anonymous):

multiply it right?

OpenStudy (jusaquikie):

yes so what is 6/6

OpenStudy (anonymous):

oh so the answer gonna 6?

OpenStudy (jusaquikie):

no, if you have a slice of pizza that is 1/6 of the pizza and then you multiply that by 6 what do you have?

OpenStudy (anonymous):